# VideoGLUE: Video General Understanding Evaluation of Foundation Models

### v1

@@ D.2 Sample-efficient transfer learning
We show our experimental results in Table 13. To better understand the data efficiency, we also show the relative Top-1 accuracy for each model (shown in the bracket), which is defined as the ratio between accuracy with fewer training examples and the accuracy achieved using all the training data. A higher relative Top-1 accuracy means the performance of the model is closer to its “full” capacity under the sample-efficient setting. We notice that the best performed model on each dataset in fully fine-tuned model also performs best in the few-shot setting. Especially, CLIP [41] achieves $46.2\%$ and $83.6\%$ relative Top-1 accuracy on K400 using only $1\%$ and $10\%$ of the training data, respectively. On SSv2, InternVideo [55] achieves $33.6\%$ and $70.6\%$ relative Top-1 accuracy with only $1\%$ and $10\%$ of the training data.

### v3

@@ E.2 Sample-efficient transfer learning
We show our experimental results in Table 14. To better understand the data efficiency, we also show the relative Top-1 accuracy for each model (shown in the bracket), which is defined as the ratio between accuracy with fewer training examples and the accuracy achieved using all the training data. A higher relative Top-1 accuracy means the performance of the model is closer to its “full” capacity under the sample-efficient setting. We notice that the best performed model on each dataset in fully fine-tuned model also performs best in the few-shot setting. Especially, CLIP (Radford et al., 2021) achieves $46.2\%$ and $83.6\%$ relative Top-1 accuracy on K400 using only $1\%$ and $10\%$ of the training data, respectively. On SSv2, InternVideo (Wang et al., 2022b) achieves $33.6\%$ and $70.6\%$ relative Top-1 accuracy with only $1\%$ and $10\%$ of the training data.